Superheated steam at 2 bar and 150oC is compressed isothermally. The steam in its final state is wet steam with a specific volume of 0.3511 m3 /kg.

Determine the change of enthalpy per kg of steam.


1
Expert's answer
2021-10-18T11:02:36-0400

"\frac mMRT_1=p_1V_1=p_2V_2,"

"p_2=\frac{RT_1} M\cdot \frac m{V_2}=3.45~bar,"

"m=\frac{p_2V_2M}{RT_1}=0.45~kg,"

"H=\frac{\gamma nR}{\gamma -1}=2.18~\frac JK ."